Réunion DUNE-IN2P3 du 7/04/21

Europe/Paris
    present - 17
Jaime Dawson, Dominique Duchesneau, Dario Autiero, Etienne Chardonnet,
Johan Collot, Alessandra Tonazzo, Andrés Munoz, Elisabetta Pennacchio,
Jean Sebastien Real, Joao Coelho, Laura Zambelli, Laurent Simard, Pablo
Kunzé, Sabrina Sacerdoti,Konstantin Shchablo, Thomas Patzak, Slavic Galymov

Laurent Vacavant, Fabien Cavalier - excused

*news -  Dario

TGIR kick off workshop will now be on remote,  25 May afternoon and
morning of 26th.

LBNC vertical drift review update discussion meetings in view of CDR
review beginning on April 28th afternoon. Then discussions on
simulation (led by task force) 10th May, and production/installation May
11th.

SPSC annual review on April 13th

------------------------------
round table

APC -
Sabrina - shows a few slides on the status of the analog optical link;
measurements of beam output from VCSEL, and the first few measurements
in liquid nitrogen. So far, two VCSELs (850nm have been tested; one
works well (better) in liquid nitrogen, the other failed (threshold
increased too high).

Etienne (analysis) continuing with thesis writing, also working on
analysis, has found a potential problem with signal deconvolution. Has
been discussing with David Adams. Slavic is surprised, thinks Gaussian
filter has already been correctly implemented.

Elisabetta asks if he will give feedback on the pre-production DP runs,
not yet however Pierre is working on it now.

--
Grenoble  
Johan Collot gave feedback from their discussions with G2ELAB related to
the HV supply. They fully understand issues but cannot suggest anything
more than the group is already aware of. Interestingly, ILL uses a 380kV
powersupply which has a 10 ppm ripple. Supply is considered to be very
good, and is very stable, had no problems of last 13 years of running.
The ripple/noise is too high, so would require filtering. These
power supplies are very expensive, 400 keuros, and still would require
filters. Grenoble (LPSC and G2ELAB) could produce the filter. Interesting
item which would be required anyway (no matter on source of HV supply).
The ILL will stop in October for 12 months, they would be able to make
tests with the power supply, possibly testing their filter.
Alternatively, buy standard 300 kV supply and add 50 kV. Worth
investigating.
In any case, filtering will probably be required. There are
commercial options for the filter, however if any modifications are
required, would be better to have experience inhouse.

Should discuss needs/requirements with Laurent and Dario.


IPNL - continuing preparations for cold box tests.

IJCLab - info given from Dominique. A modification was needed to make the
chimney test in the IJCLab cryostat.


LAPP - Dominique
working on method to assemble CRP. Anode design finalised last week,
sent to ELTOS. Now working on how to assemble.
Design the composite structure which will hold anodes is being worked on.
May be shown next week.

analyse - Laura has discussed with Filippo for space charge simulation
(COMSOL).
- Pablo, working on track selection, E field correction
There are minutes attached to this event. Show them.
    • 10:00 10:10
      Nouvelles récentes du projet 10m
      Speaker: Dario Autiero (IPNL)
    • 10:10 10:25
      Quelques 'highlights' côté activités des labos 15m
      Speaker: Tour de table
    • 10:25 10:40
      Chapitre choisi côté technique (TBD) 15m
    • 10:40 10:55
      Morceaux choisis sur analyse: tbd 15m
    • 10:55 11:00
      A.O.B. 5m